# Jacquet modules of Tate cohomology and base change lifting

### v1

@@ 6.1. Mirabolic restriction of $\theta_{10}$
for all $\varphi\in\overline{\mathbb{Q}}_{\ell}[\mathbb{F}_{q}]$. Let $A(\mathbb{F}_{q})$ be the subgroup {±$\{\pm I_{4}\}\,{\rm SL}_{2}(\mathbb{F}_{q})H(\mathbb{F}_{q})$, where $I_{4}$ is the $4\times 4$ identity matrix. The $H(\mathbb{F}_{q})$ representation $\xi$ is extended to a representation of $A(\mathbb{F}_{q})$ where the actions of {±$\{\pm I_{4}\}$ and ${\rm SL}_{2}(\mathbb{F}_{q})$ are defined as in [Gol78, equations $(1)$ and $(2)$]. Then the restriction ${\rm res}_{{\rm SL}_{2}(\mathbb{F}_{q})}(\xi)$ splits into two irreducible representations $\xi^{+}$ and $\xi^{-}$ acting on the spaces of even and odd functions in ¯$\overline{\mathbb{Q}}_{\ell}[\mathbb{F}_{q}]$, respectively. We further extend the representation $\xi^{-}$ trivially to the group $A(\mathbb{F}_{q})$ and set $\widetilde{\xi}=\xi\otimes\xi^{-}$. The induced representation ${\rm Ind}_{A(\mathbb{F}_{q})}^{Q(\mathbb{F}_{q})}(\widetilde{\xi})$ is irreducible, which we denote by $\pi_{1}$. Similarly, we have an irreducible representation $\pi_{1,\tau}={\rm Ind}_{A(\mathbb{F}_{q})}^{Q(\mathbb{F}_{q})}(\widetilde{\xi_{\tau}})$, where the representation $\widetilde{\xi_{\tau}}$ is defined analogously to $\widetilde{\xi}$ by replacing the character $\psi(t)$ by $\psi_{\tau}(t)=\psi(\tau t)$. Then we have the isomorphism of $Q(\mathbb{F}_{q})$ representations ([Gol78, Proposition 2])
${\rm res}_{Q(\mathbb{F}_{q})}(\pi)\simeq\pi_{1}\oplus\pi_{1,\tau}.$
6.3.2.
Using similar constructions as in the preceding subsection, we get that the restriction ${\rm res}_{Q(\mathbb{F}_{q^{\ell}})}(\Pi)$ is isomorphic to the direct sum $\Pi_{1}\oplus\Pi_{1,\tau}$, where $\Pi_{1}$ and $\Pi_{1,\alpha}$ are the irreducible induced representations ${\rm Ind}_{A(\mathbb{F}_{q^{\ell}})}^{Q(\mathbb{F}_{q^{\ell}})}(\widetilde{\xi}_{\ell})$ and ${\rm Ind}_{A(\mathbb{F}_{q^{\ell}})}^{Q(\mathbb{F}_{q^{\ell}})}(\widetilde{\xi_{\tau}}_{\ell})$ respectively, and the irreducible representations $\widetilde{\xi}_{\ell}$ (resp. $\widetilde{\xi_{\tau}}_{\ell}$) is defined analogously to $\widetilde{\xi}$ (resp. $\widetilde{\xi_{\tau}}$) by replacing the character $\psi$ by $\eta=\psi\circ{\rm Tr}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}$. Note that $\Pi_{1}$ and $\Pi_{1,\tau}$ are stable under the action of $\langle\sigma\rangle$, given by
$(\sigma f)(t)=f(\sigma^{-1}(t)),\,\,t\in\mathbb{F}_{q^{\ell}},\,f\in\overline{\mathbb{Q}}_{\ell}[\mathbb{F}_{q^{\ell}}].$
Let $\chi_{\widetilde{\xi}_{\ell}}$ and $\chi_{\widetilde{\xi_{\tau}}_{\ell}}$ be the characters of $\widetilde{\xi}_{\ell}$ and $\widetilde{\xi_{\tau}}_{\ell}$ respectively. Then using the relation (6.2) and [Gol78, Equations $(1)$ and $(2)$], we get the following identity of characters:
@@
for all $g\in A(\mathbb{F}_{q})$. Using these, we now deduce the following theorem.
Theorem 6.3.
Let $\pi$ and $\Pi$ be the unique irreducible unipotent cuspidal $\ell$-adic representation of ${\rm Sp}_{4}(\mathbb{F}_{q})$ and ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})$, respectively. Then the Tate cohomology group $\widehat{H}^{0}(r_{\ell}(\Pi))$ is isomorphic to the Frobenius twist $r_{\ell}(\pi)^{(\ell)}$.
@@
for all $g\in{\rm Sp}_{4}(\mathbb{F}_{q})$. Let $g=su=us$ be the Jordan decomposition of $g$. Note that either $g$ is contained in a parabolic subgroup or $u$ is the identity element and $s$ is regular. If $s$ is a regular semisimple element, we have $\chi_{\Pi}(s)=1=\chi_{\pi}(s)$. Assume that $g$ belongs to a parabolic subgroup $P(\mathbb{F}_{q})$ or $Q(\mathbb{F}_{q})$. If $g\in P(\mathbb{F}_{q})$, then it follows from the arguments of Lemma 6.2 that
@@
Suppose $g\in Q(\mathbb{F}_{q})$. Note that $\chi_{\Pi}=\chi_{\Pi_{1}}+\chi_{\Pi_{1,\tau}}$ and $\chi_{\pi}=\chi_{\pi_{1}}+\chi_{\pi_{1,\tau}}$. Now, using the character formula for induced representations, we have
$\chi_{\Pi_{1}}(g)=\sum_{y\in Y_{\ell},\,ygy^{-1}\in A(\mathbb{F}_{q^{\ell}})}\chi_{\widetilde{\xi}_{l}}(ygy^{-1})$ (6.4)
@@
where $Y_{\ell}$ and $Y$ denote the right coset spaces $A(\mathbb{F}_{q^{\ell}})\backslash Q(\mathbb{F}_{q^{\ell}})$ and $A(\mathbb{F}_{q})\backslash Q(\mathbb{F}_{q})$ respectively, with $Y_{\ell}^{\sigma}=Y$. Since $\langle\sigma\rangle$ acts freely on $Y_{\ell}\setminus Y$, there exists a set $\mathcal{U}\subseteq Y_{\ell}\setminus Y$ such that $Y_{\ell}\setminus Y$ is the disjoint union of $\sigma^{i}\mathcal{U}$, $10\leq i\leq\ell-$ 1. Then the identity (6.4) becomes
@@
Using the identity (6.3) and taking $-\ell$ reduction to the above relation, we get
@@
for all $g\in Q(\mathbb{F}_{q})$. This completes the proof.

### v2

@@ 7.1. Mirabolic restriction of $\theta_{10}$
for $h\in H(\mathbb{F}_{q})$ and $\varphi\in\overline{\mathbb{Q}}_{\ell}[\mathbb{F}_{q}]$. Let $A(\mathbb{F}_{q})$ be the subgroup {±$\{\pm I_{4}\}\,{\rm SL}_{2}(\mathbb{F}_{q})H(\mathbb{F}_{q})$, where $I_{4}$ is the $\times 4$ 4 identity matrix. The $H(\mathbb{F}_{q})$ representation $\xi$ is extended to a representation of $A(\mathbb{F}_{q})$ where the actions of {±$\{\pm I_{4}\}$ and ${\rm SL}_{2}(\mathbb{F}_{q})$ are defined as in [Gol78, equations (1)(1)( 1 ) and (2)(2)( 2 )]. The restriction ${\rm res}_{{\rm SL}_{2}(\mathbb{F}_{q})}(\xi)$ splits into two irreducible representations $\xi^{+}$ and $\xi^{-}$ acting on the spaces of even and odd functions in ¯$\overline{\mathbb{Q}}_{\ell}[\mathbb{F}_{q}]$, respectively. We further extend the representation $\xi^{-}$ trivially to the group $A(\mathbb{F}_{q})$ and set $\widetilde{\xi}=\xi\otimes\xi^{-}$. The induced representation ${\rm Ind}_{A(\mathbb{F}_{q})}^{Q(\mathbb{F}_{q})}(\widetilde{\xi})$ is irreducible, which we denote by $\pi_{1}$. Similarly, we have an irreducible representation $\pi_{1,\tau}={\rm Ind}_{A(\mathbb{F}_{q})}^{Q(\mathbb{F}_{q})}(\widetilde{\xi_{\tau}})$, where $\widetilde{\xi_{\tau}}$ is the representation of $A(\mathbb{F}_{q})$ defined analogously to $\widetilde{\xi}$ by replacing the character $\psi(t)$ by $\psi_{\tau}(t)=\psi(\tau t)$. Then the restriction ${\rm res}_{Q(\mathbb{F}_{q})}(\pi)$ is isomorphic to the direct sum $\pi_{1}\oplus\pi_{1,\tau}$ ([Gol78, Proposition 2]).
7.3.2.
Using similar constructions as in the preceding subsection, we get that the restriction ${\rm res}_{Q(\mathbb{F}_{q^{\ell}})}(\Pi)$ is isomorphic to the direct sum $\Pi_{1}\oplus\Pi_{1,\tau}$, where $\Pi_{1}$ and $\Pi_{1,\alpha}$ are the irreducible induced representations ${\rm Ind}_{A(\mathbb{F}_{q^{\ell}})}^{Q(\mathbb{F}_{q^{\ell}})}(\widetilde{\xi}_{\ell})$ and ${\rm Ind}_{A(\mathbb{F}_{q^{\ell}})}^{Q(\mathbb{F}_{q^{\ell}})}(\widetilde{\xi_{\tau}}_{\ell})$ respectively, and the irreducible representations $\widetilde{\xi}_{\ell}$ (resp. $\widetilde{\xi_{\tau}}_{\ell}$) is defined analogously to $\widetilde{\xi}$ (resp. $\widetilde{\xi_{\tau}}$) by replacing the character $\psi$ by $\eta=\psi\circ{\rm Tr}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}$. Note that $\Pi_{1}$ and $\Pi_{1,\tau}$ are endowed with an action of $\Gamma$, given by
$(\sigma f)(x)=f(\sigma^{-1}(x)),\,\,\forall x\in Q(\mathbb{F}_{q^{\ell}}),\,f\in\Pi_{1}\,(\text{or}\,\,\Pi_{1,\tau}).$
Let $\chi_{\widetilde{\xi}_{\ell}}$ and $\chi_{\widetilde{\xi_{\tau}}_{\ell}}$ be the characters of $\widetilde{\xi}_{\ell}$ and $\widetilde{\xi_{\tau}}_{\ell}$ respectively. Then using the relation (7.2) and [Gol78, Equations (1)(1)( 1 ) and (2)(2)( 2 )], we get the following identity of characters:
@@
for all $g\in A(\mathbb{F}_{q})$. We now prove the following theorem.
Theorem 7.3.1.
Let $\ell$ be an odd prime number. Let $\Gamma$ be the Galois group ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$, and let $\sigma$ be the Frobenius element in $\Gamma$. Let $\pi$ and $\Pi$ be the irreducible unipotent $\ell$-adic cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q})$ and ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})$, respectively. Then the Tate cohomology group $\widehat{H}^{0}(r_{\ell}(\Pi))$ is isomorphic to the Frobenius twist $r_{\ell}(\pi)^{(\ell)}$.
@@
for all $g\in{\rm Sp}_{4}(\mathbb{F}_{q})$. Let $g=su=us$ be the Jordan decomposition of $g$. Note that either $g$ is contained in a parabolic subgroup of ${\rm Sp}_{4}(\mathbb{F}_{q})$ or $u$ is the identity element and $s$ is regular. If $s$ is a regular semisimple element, we have $\chi_{\Pi}(s)=1=\chi_{\pi}(s)$. Assume that $g$ belongs to a parabolic subgroup $P(\mathbb{F}_{q})$ or $Q(\mathbb{F}_{q})$. If $g\in P(\mathbb{F}_{q})$, then it follows from the arguments of Lemma 7.2.2 that
@@
Suppose $g\in Q(\mathbb{F}_{q})$. Note that $\chi_{\Pi}=\chi_{\Pi_{1}}+\chi_{\Pi_{1,\tau}}$ and $\chi_{\pi}=\chi_{\pi_{1}}+\chi_{\pi_{1,\tau}}$. Using character formula for induced representations, we have
$\chi_{\Pi_{1}}(g)=\sum_{y\in Y_{\ell},\,ygy^{-1}\in A(\mathbb{F}_{q^{\ell}})}\chi_{\widetilde{\xi}_{\ell}}(ygy^{-1})$ (7.4)
@@
where $Y_{\ell}$ and $Y$ denote the right coset spaces $A(\mathbb{F}_{q^{\ell}})\backslash Q(\mathbb{F}_{q^{\ell}})$ and $A(\mathbb{F}_{q})\backslash Q(\mathbb{F}_{q})$ respectively, with $Y_{\ell}^{\sigma}=Y$. Since $\Gamma$ acts freely on $Y_{\ell}\setminus Y$, there exists a set $\mathcal{U}\subseteq Y_{\ell}\setminus Y$ such that $Y_{\ell}\setminus Y$ is the disjoint union of $\sigma^{i}\mathcal{U}$, $0\leq i\leq\ell-1$. Also note that $\chi_{\widetilde{\xi}_{\ell}}(\sigma(a))=\chi_{\widetilde{\xi}_{\ell}}(a)$ for all $a\in\mathbb{A}(\mathbb{F}_{q^{\ell}})$. Then the identity (7.4) becomes
@@
Using the relations (7.3) and taking $-\ell$ reduction to the above equation, we get
@@
for all $g\in{\rm Sp}_{4}(\mathbb{F}_{q})$. Using [Lam01, Corollary 7.21], we conclude that the Tate cohomology $\widehat{H}^{0}(r_{\ell}(\Pi))$ is isomorphic to $r_{\ell}(\pi)^{(\ell)}$.
8. Conflict of interest statement
On behalf of all authors, the corresponding author states that there is no conflict of interest.
9. Date statement
No associated data